### v2

# Results related to the Gaussian product inequality conjecture for mixed-sign exponents in arbitrary dimension

Abstract
This paper establishes that the opposite Gaussian product inequality (GPI) of the type proved by Russell & Sun, (2022a) in two dimensions, and partially extended to higher dimensions by Zhou et al. , (2024), continues to hold for an arbitrary mix of positive and negative exponents. A general quantitative lower bound is also obtained conditionally on the GPI conjecture being true.
keywords:
centered Gaussian random vector , convex order , Gaussian product inequality , Loewner order , multivariate normal distribution , opposite Gaussian product inequality , real linear polarization constant
MSC:
[2020]Primary: 60E15; Secondary 26A48, 44A10, 62E15, 62H10, 62H12
journal: Statistics & Probability Letters
1 Introduction
The real linear polarization constant conjecture grew out of the work of Benítez et al. , (1998) on obtaining sharp lower bounds for the norms of products of polynomials on real Banach spaces, and, in the linear case, products of bounded linear functionals. A decade later, Frenkel, (2008) recast the problem restricted to linear functionals on a real Euclidean space as the following Gaussian product inequality (GPI) conjecture: for every real centered Gaussian random vector $\boldsymbol{X}=(X_{1},\ldots,X_{d})$ with $d\in\mathbb{N}$ and each $m\in\mathbb{N}$,
$\mathsf{E}\left[\prod_{i=1}^{d}X_{i}^{2m}\right]\geq\prod_{i=1}^{d}\mathsf{E}\left[X_{i}^{2m}\right].$ (1)
Inequality (1) implies the (restricted) real linear polarization constant conjecture (Malicet et al. ,, 2016). It is also closely linked to the $U$-conjecture: if two polynomials in the components of $\boldsymbol{X}$ are independent, an orthogonal transformation can be found so that each depends on a disjoint subset of variables; see, e.g., Kagan et al. , (1973); Malicet et al. , (2016).
Building on this formulation, Li & Wei, (2012) proposed a strengthened conjecture allowing arbitrary positive exponents, represented as $\boldsymbol{\nu}=(\nu_{1},\ldots,\nu_{d})\in(0,\infty)^{d}$. The statement, denoted by $\mathbf{GPI}_{d}(\boldsymbol{\nu})$, reads
$\mathbf{GPI}_{d}(\boldsymbol{\nu}):\quad\mathsf{E}\left[\prod_{i=1}^{d}|X_{i}|^{2\nu_{i}}\right]\geq\prod_{i=1}^{d}\mathsf{E}\left[|X_{i}|^{2\nu_{i}}\right].$ (2)
The conjecture is settled for $d=2$ because $(|X_{1}|,|X_{2}|)$ is multivariate totally positive of order 2 (Karlin & Rinott,, 1981). In dimension $3$, numerous special cases have been confirmed:
  1. (a)
    Lan et al. , (2020): $\mathbf{GPI}_{3}(p,p,q)$ for all $p,q\in\mathbb{N}$;
  2. (b)
    Russell & Sun, (2023): $\mathbf{GPI}_{3}(1,p,q)$ for all $p,q\in\mathbb{N}$;
  3. (c)
    Russell & Sun, (2024): $\mathbf{GPI}_{3}(2,3,p)$ for all $p\in\mathbb{N}$;
  4. (d)
    Herry et al. , (2024): $\mathbf{GPI}_{3}(p,q,r)$ for all $p,q,r\in\mathbb{N}$;
  5. (e)
    Kim & Kim, (2025): $\mathbf{GPI}_{3}(p,q,r)$ for $p\in\mathbb{N}$ and $q,r\in(0,\infty)$;
  6. (f)
    Kim et al. , (2025): $\mathbf{GPI}_{3}(p,q,r)$ for all $p,q,r\in(1/2)\mathbb{N}$.
For higher $d$, additional results appear under various covariance constraints; see, for instance, Genest & Ouimet, (2022), Russell & Sun, (2022b), and Edelmann et al. , (2023).
Interest has also turned to scenarios where some exponents in (2) are negative (subject to the expectations being finite). When every exponent is negative, Wei, (2014) proved the inequality and even extended it to multivariate gamma random vectors (also called permanental vectors). Wei’s results, and related developments, extend to traces and determinants of disjoint principal blocks of Wishart random matrices, as shown by Genest & Ouimet, (2023); Genest et al. , (2024, 2025).
In the Gaussian setting, the mixed-sign case, where both positive and negative exponents appear, remains largely unresolved. The two-dimensional situation was settled by Russell & Sun, (2022a) (tighter quantitative bounds have also been proved by Hu et al. , (2023)), and partial progress for general $d$ was achieved recently by Zhou et al. , (2024). Letting $\Sigma=(\sigma_{ij})_{1\leq i,j\leq d}$ be the covariance matrix of $\boldsymbol{X}$, they established
$\mathsf{E}\left[|X_{1}|^{-2\nu_{1}}\prod_{i=2}^{d}X_{i}^{2}\right]\geq\left\{\prod_{i=2}^{d}\left[1-\frac{\sigma_{1i}^{2}}{\sigma_{11}\sigma_{ii}}\right]\right\}\mathsf{E}\left[|X_{1}|^{-2\nu_{1}}\right]\prod_{i=2}^{d}\mathsf{E}[X_{i}^{2}],$ (3)
for any $\nu_{1}\in[0,1/2)$, and
$\mathsf{E}\left[\left(\prod_{i=1}^{d-1}|X_{i}|^{-2\nu_{i}}\right)|X_{d}|^{2\nu_{d}}\right]\leq\mathsf{E}\left[\prod_{i=1}^{d-1}|X_{i}|^{-2\nu_{i}}\right]\mathsf{E}\left[|X_{d}|^{2\nu_{d}}\right],$ (4)
for $\nu_{1},\ldots,\nu_{d-1}\in[0,1/2)$ and $\nu_{d}\in(0,\infty)$.
The present paper investigates the general mixed-sign problem, thereby significantly generalizing (3) and (4). Except for a narrow unresolved case discussed in Remark 2, the analysis resolves the question through Theorems 1 and 2.
The remainder of the paper is organized as follows. Section 2 states the main results; proofs are given in Section 3. In the interest of self-containment, some technical lemmas used in the proofs are relegated to Appendix A.
2 Results
The first theorem below generalizes (3) from one Gaussian component having a negative exponent to an arbitrary subset of the Gaussian components having negative exponents. The positive exponents $\nu_{i}$ are not restricted to be equal to $1$, but this comes at the cost of assuming the validity of a lower-order GPI. However, this also means that if the GPI assumption is known to hold in the literature, then the result of the theorem holds ‘unconditionally’. For a comprehensive list of results related to the GPI, refer to Section 1 of Genest et al. , (2024).
Theorem 1 (Lower bound).
Let $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma)$ for some positive definite matrix $\Sigma$, and let $\emptyset\neq\mathcal{J}\subseteq\{1,\ldots,d\}$ be given. Assume that $\nu_{j}\in[0,1/2)$ for all $j\in\mathcal{J}$ and $\nu_{i}\in(0,\infty)$ for all $i\in\mathcal{J}^{\complement}$. Then, conditionally on $\mathbf{GPI}_{d-|\mathcal{J}|}((\nu_{i})_{i\in\mathcal{J}^{\complement}})$ being true, we have
$\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\times\prod_{i\in\mathcal{J}^{\complement}}|X_{i}|^{2\nu_{i}}\right]\geq\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\right]\times\prod_{i\in\mathcal{J}^{\complement}}\left\{\frac{(\Sigma/\Sigma_{\mathcal{J}\mathcal{J}})_{ii}}{\Sigma_{ii}}\right\}^{\nu_{i}}\mathsf{E}\left[|X_{i}|^{2\nu_{i}}\right],$
where $\mathcal{J}^{\complement}\equiv\{1,\ldots,d\}\backslash\mathcal{J}$, the matrix $\Sigma_{\mathcal{I}\mathcal{I}^{\prime}}$ denotes $\Sigma$ restricted to the rows and columns indexed in $\mathcal{I}$ and $\mathcal{I}^{\prime}$, respectively, and $\Sigma/\Sigma_{\mathcal{J}\mathcal{J}}=\Sigma_{\mathcal{J}^{\complement}\mathcal{J}^{\complement}}-\Sigma_{\mathcal{J}^{\complement}\mathcal{J}}\Sigma_{\mathcal{J}\mathcal{J}}^{-1}\Sigma_{\mathcal{J}\mathcal{J}^{\complement}}$ is the Schur complement of $\Sigma_{\mathcal{J}\mathcal{J}}$ in $\Sigma$.
Remark 1.
In Theorem 1, the factor $(\Sigma/\Sigma_{\mathcal{J}\mathcal{J}})_{ii}/\Sigma_{ii}$ is smaller or equal to $1$ because $\Sigma/\Sigma_{\mathcal{J}\mathcal{J}}\preceq\Sigma$ in the Loewner order. In the special case where $\mathcal{J}=\{1\}$ and $\mathbf{GPI}_{d}(1,\ldots,1)$ is known to hold (Frenkel,, 2008), then we recover (3) with
$\frac{(\Sigma/\Sigma_{\mathcal{J}\mathcal{J}})_{ii}}{\Sigma_{ii}}=\frac{\sigma_{ii}-\sigma_{i1}\sigma_{11}^{-1}\sigma_{1i}}{\sigma_{ii}}=1-\frac{\sigma_{1i}^{2}}{\sigma_{11}\sigma_{ii}},\quad i\in\{2,\ldots,d\}.$
From the work of Wei, (2014), the GPI is known to hold when all the exponents are negative, so the following corollary is immediate.
Corollary 1.
In the setting of Theorem 1, we have
$\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\times\prod_{i\in\mathcal{J}^{\complement}}|X_{i}|^{2\nu_{i}}\right]\geq\prod_{j\in\mathcal{J}}\mathsf{E}\left[|X_{j}|^{-2\nu_{j}}\right]\times\prod_{i\in\mathcal{J}^{\complement}}\left\{\frac{(\Sigma/\Sigma_{\mathcal{J}\mathcal{J}})_{ii}}{\Sigma_{ii}}\right\}^{\nu_{i}}\mathsf{E}\left[|X_{i}|^{2\nu_{i}}\right].$
Next, the goal is to generalize (4) by replacing $|X_{d}|^{2\nu_{d}}$ with a product of positive powers of the absolute Gaussian components indexed in $\mathcal{J}^{\complement}$ or, more generally, a convex function of such components.
Theorem 2 (Upper bound).
Let $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma)$ for some positive definite matrix $\Sigma$, and let $\emptyset\neq\mathcal{J}\subseteq\{1,\ldots,d\}$ be given. Assume that $\nu_{j}\in[0,1/2)$ for all $j\in\mathcal{J}$. Then, for any convex function $\psi:\mathbb{R}^{d-|\mathcal{J}|}\to\mathbb{R}$, we have
$\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\times\psi((X_{i})_{i\in\mathcal{J}^{\complement}})\right]\leq\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\right]\times\mathsf{E}\left[\psi((X_{i})_{i\in\mathcal{J}^{\complement}})\right].$
Since $(x_{i})_{i\in\mathcal{J}^{\complement}}\mapsto(\sum_{i\in\mathcal{J}^{\complement}}\frac{\nu_{i}}{\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}}|x_{i}|)^{2\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}}$ is convex for $\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}\in[1/2,\infty)$, the following corollary is immediate by combining Theorem 2 with the weighted arithmetic mean-geometric mean inequality.
Corollary 2.
Let $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma)$ for some positive definite matrix $\Sigma$, and let $\emptyset\neq\mathcal{J}\subseteq\{1,\ldots,d\}$ be given. Assume that $\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}\in[1/2,\infty)$. Then, we have
$\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\times\prod_{i\in\mathcal{J}^{\complement}}|X_{i}|^{2\nu_{i}}\right]\leq\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\right]\times\mathsf{E}\left[\left(\sum_{i\in\mathcal{J}^{\complement}}\frac{\nu_{i}}{\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}}|X_{i}|\right)^{2\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}}\right].$
Remark 2.
The case $\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}\in(0,1/2)$ remains open. One possible approach would be to exploit the Bernstein function integral representation, $|x|^{2\nu}=\nu\int_{0}^{\infty}(1-e^{-sx^{2}})s^{-\nu-1}\mathrm{d}s/\Gamma(1-\nu)$, which holds for $x\in\mathbb{R}$ and $\nu\in(0,1/2)$; see, e.g., Schilling et al. , (2012, Eq. (1)). However, we were unable to turn this representation into a successful argument. A treatment of this case therefore lies beyond the scope of our paper.
In the two-dimensional case, we have the following generalization of Corollary 2.
Proposition 1.
Suppose that the two-dimensional random vector $(X,Y)$ has an elliptical distribution, meaning that $(X,Y)\smash{\stackrel{{\scriptstyle\mathrm{law}}}{{=}}}R\Sigma^{1/2}\boldsymbol{U}$, where $R$ is a nonnegative random variable (the radial part), $\Sigma^{1/2}$ is the symmetric square root of some positive definite matrix $\Sigma$, and $\boldsymbol{U}$ is a random vector uniformly distributed on the unit circle $\mathbb{S}^{1}$ that is independent of $R$. Also, let $f$ and $g$ be any nonincreasing and nondecreasing functions on $[0,\infty)$, respectively. Then
$\mathsf{E}[f(|X|)g(|Y|)]\leq\mathsf{E}[f(|X|)]\mathsf{E}[g(|Y|)].$
3 Proofs
Proof of Theorem 1.
Without loss of generality, assume that $\nu_{j}\in(0,1/2)$ for all $j\in\mathcal{J}$, and re-index the components of $\boldsymbol{X}$ so that $\mathcal{J}=\{1,\ldots,k\}$ with $11\leq k\leq d-$ 1 and $\mathcal{J}^{\complement}=\{k+1,\ldots,d\}$. For every $x\in\mathbb{R}$ and $\nu\in(0,1/2)$, the identity
$|x|^{-2\nu}=\frac{1}{\Gamma(\nu)}\int_{0}^{\infty}e^{-sx^{2}}s^{\nu-1}\mathrm{d}s$ (5)
holds; see, e.g., Schilling et al. , (2012, Eq. (2)). Applying (5) to $|X_{1}|^{-2\nu_{1}},\ldots,|X_{k}|^{-2\nu_{k}}$ and using Fubini’s theorem yields
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]$ (6)
$\qquad=\frac{1}{\prod_{j=1}^{k}\Gamma(\nu_{j})}\int_{(0,\infty)^{k}}\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]\prod_{j=1}^{k}s_{j}^{\nu_{j}-1}\mathrm{d}\boldsymbol{s},$
where $\boldsymbol{s}=(s_{1},\ldots,s_{k})$ and $T_{\boldsymbol{s}}=\mathrm{diag}(2s_{1},\ldots,2s_{k},0,\ldots,0)$. By a renormalization argument, we have
$\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]$
$\qquad=\int_{(0,\infty)^{d}}\exp(-\boldsymbol{x}^{\top}T_{\boldsymbol{s}}\boldsymbol{x}/2)\,\frac{\exp(-\boldsymbol{x}^{\top}\Sigma^{-1}\boldsymbol{x}/2)}{(2\pi)^{d/2}|\Sigma|^{1/2}}\prod_{i=k+1}^{d}|x_{i}|^{2\nu_{i}}\mathrm{d}\boldsymbol{x}$
$\qquad=\frac{|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}{|\Sigma|^{1/2}}\int_{(0,\infty)^{d}}\frac{\exp\{-\boldsymbol{x}^{\top}(\Sigma^{-1}+T_{\boldsymbol{s}})\,\boldsymbol{x}/2\}}{(2\pi)^{d/2}|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}\prod_{i=k+1}^{d}|x_{i}|^{2\nu_{i}}\mathrm{d}\boldsymbol{x}$
$\qquad=\frac{|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}{|\Sigma|^{1/2}}\,\mathsf{E}\left[\prod_{i=k+1}^{d}|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right],$
where $\smash{\boldsymbol{Y}^{(\boldsymbol{s})}}=(Y_{1}^{(\boldsymbol{s})},\ldots,Y_{d}^{(\boldsymbol{s})})\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1})$. Then, using the assumption that $\mathbf{GPI}_{d-k}((\nu_{i})_{i>k})$ holds, we obtain
$\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]$ $\geq\frac{|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}{|\Sigma|^{1/2}}\,\prod_{i=k+1}^{d}\mathsf{E}\left[|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right]$
$=\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\right]\,\prod_{i=k+1}^{d}\mathsf{E}\left[|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right].$
By applying the last bound in (6) and using (5) together with Fubini’s theorem, it follows that
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]\geq\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\right]\prod_{i=k+1}^{d}\mathsf{E}\left[|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right].$ (7)
Finally, to evaluate $\mathsf{E}\left[|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right]$, note that, for $i>k$,
$\mathsf{E}\left[|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right]$ $=2\int_{0}^{\infty}u^{2\nu_{i}}\frac{1}{\sqrt{2\pi\,\mathsf{Var}(Y_{i}^{(\boldsymbol{s})})}}\exp\left\{-\frac{u^{2}}{2\,\mathsf{Var}(Y_{i}^{(\boldsymbol{s})})}\right\}\mathrm{d}u$
$=\frac{2^{\nu_{i}}\{\mathsf{Var}(Y_{i}^{(\boldsymbol{s})})\}^{\nu_{i}}}{\sqrt{\pi}}\int_{0}^{\infty}t^{\nu_{i}-1/2}e^{-t}\mathrm{d}t=\frac{\Gamma(\nu_{i}+1/2)}{\sqrt{\pi}}2^{\nu_{i}}\{\mathsf{Var}(Y_{i}^{(\boldsymbol{s})})\}^{\nu_{i}},$
by applying the change of variable $t=u^{2}/(2\,\mathsf{Var}(Y_{i}^{(\boldsymbol{s})}))$. Next, write $\Sigma$ in block form:
$\Sigma=\begin{bmatrix}A&B\\ B^{\top}&C\end{bmatrix},$
where $A$ is $k\times k$, $B$ is $k\times(d-k)$ and $C$ is $(d-k)\times(d-k)$. Let $S=\mathrm{diag}(2s_{1},\ldots,2s_{k})$ so that $T_{\boldsymbol{s}}=\mathrm{diag}(S,0_{(d-k)\times(d-k)})$. The matrix inversion formula for block partitions in Lemma 1 (i) gives
$\Sigma^{-1}+T_{\boldsymbol{s}}=\begin{bmatrix}A^{-1}+S+A^{-1}B(\Sigma/A)^{-1}B^{\top}A^{-1}&-A^{-1}B(\Sigma/A)^{-1}\\[2.84526pt] -(\Sigma/A)^{-1}B^{\top}A^{-1}&(\Sigma/A)^{-1}\end{bmatrix}\equiv\begin{bmatrix}P&Q\\ Q^{\top}&R\end{bmatrix}.$
Then Lemma 1 (ii) yields
$(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}$ $=\begin{bmatrix}(A^{-1}+S)^{-1}&\star\\[2.84526pt] \star&(\Sigma/A)+B^{\top}A^{-1}(A^{-1}+S)^{-1}A^{-1}B\end{bmatrix}.$
Let $\boldsymbol{e}_{j}$ denote the $j$th standard basis vector. Since $B^{\top}A^{-1}(A^{-1}+S)^{-1}A^{-1}B$ is positive definite and $\mathsf{Var}(X_{i})=\Sigma_{ii}$, then, for $i>k$,
$\mathsf{Var}(Y^{(\boldsymbol{s})}_{i})$ $=\boldsymbol{e}_{i-k}^{\top}\{(\Sigma/A)+B^{\top}A^{-1}(A^{-1}+S)^{-1}A^{-1}B\}\boldsymbol{e}_{i-k}$
$\geq\boldsymbol{e}_{i-k}^{\top}(\Sigma/A)\boldsymbol{e}_{i-k}$
$=\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\mathsf{Var}(X_{i}),$
and therefore,
$\mathsf{E}\left[|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right]$$\geq\left\{\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\right\}^{\nu_{i}}\frac{\Gamma(\nu_{i}+1/2)}{\sqrt{\pi}}2^{\nu_{i}}\{\mathsf{Var}(X_{i})\}^{\nu_{i}}$
=$=\left\{\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\right\}^{\nu_{i}}\mathsf{E}[|X_{i}|^{2\nu_{i}}].$
The conclusion follows by plugging this lower bound in (7). ∎
Proof of Theorem 2.
Following the same steps as in the proof of Theorem 1 but without applying $\mathbf{GPI}_{d-k}((\nu_{i})_{i>k})$, we have
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\times\psi((X_{i})_{i>k})\right]$
$\qquad=\frac{1}{\prod_{j=1}^{k}\Gamma(\nu_{j})}\int_{(0,\infty)^{k}}\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\psi((X_{i})_{i>k})\right]\prod_{j=1}^{k}s_{j}^{\nu_{j}-1}\mathrm{d}\boldsymbol{s},$
with
$\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\psi((X_{i})_{i>k})\right]=\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\right]\,\mathsf{E}\left[\psi((Y_{i}^{(\boldsymbol{s})})_{i>k})\right].$
Since $\Sigma^{-1}\preceq\Sigma^{-1}+T_{\boldsymbol{s}}$ in the Loewner order, then $(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}\preceq\Sigma$, and thus $\smash{\boldsymbol{Y}^{(\boldsymbol{s})}}$ is smaller or equal to $\boldsymbol{X}$ in the convex order, written $\smash{\boldsymbol{Y}^{(\boldsymbol{s})}}\leq_{\mathrm{cx}}\boldsymbol{X}$, by Lemma 2, so that
$\mathsf{E}\left[\psi((Y_{i}^{(\boldsymbol{s})})_{i>k})\right]\leq\mathsf{E}\left[\psi((X_{i})_{i>k})\right].$
Combining the last two equations and applying (5) again together with Fubini’s theorem, we find that
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\times\psi((X_{i})_{i>k})\right]\leq\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\right]\times\mathsf{E}\left[\psi((X_{i})_{i>k})\right].$
This concludes the proof. ∎
Proof of Proposition 1.
Let $f$ and $g$ be any nonincreasing and nondecreasing functions on $[0,\infty)$, respectively. For all $r\in[0,\infty)$, denote the scaled functions $f_{r}(x)=f(rx)$ and $g_{r}(y)=g(ry)$. Given the stochastic representation of elliptical distributions in the statement of the proposition, and the rotational invariance of $\boldsymbol{U}=(U_{1},U_{2})$, it is sufficient to show that, for all $r\in[0,\infty)$,
$\mathsf{E}[f_{r}(|(\Sigma^{1/2}\boldsymbol{U})_{1}|)g_{r}(|(\Sigma^{1/2}\boldsymbol{U})_{2}|)]\leq\mathsf{E}[f_{r}(|U_{1}|)]\,\mathsf{E}[g_{r}(|U_{2}|)],$
because we can just integrate the above inequality on both sides with respect to the distribution of $R$ to obtain the inequality claimed in the statement of the proposition. Moreover, for $r\in[0,\infty)$, the scaled functions $f_{r}$ and $g_{r}$ remain nonincreasing and nondecreasing, respectively. Therefore, to conclude, we can ignore $r$ and any scalings related to the rows of $\Sigma^{1/2}$; it is sufficient to prove
$\mathsf{E}[f(|\boldsymbol{\alpha}\cdot\boldsymbol{U}|)g(|\boldsymbol{\beta}\cdot\boldsymbol{U}|)]\leq\mathsf{E}[f(|\boldsymbol{e}_{1}\cdot\boldsymbol{U}|)]\,\mathsf{E}[g(|\boldsymbol{e}_{2}\cdot\boldsymbol{U}|)],$ (8)
where $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ denote the first and second rows of $\Sigma^{1/2}$ under the assumption $\|\boldsymbol{\alpha}\|_{2}=\|\boldsymbol{\beta}\|_{2}=1$, and $\boldsymbol{e}_{i}$ denotes the $i$th standard basis vector.
To this end, consider $T$ the $2\times 2$ matrix for a $\pi/2$ counterclockwise rotation, and the function
$H(\boldsymbol{u})=\{f(|\boldsymbol{\alpha}\cdot\boldsymbol{u}|)-f(|T\boldsymbol{\beta}\cdot\boldsymbol{u}|)\}\{g(|\boldsymbol{\beta}\cdot\boldsymbol{u}|)-g(|T\boldsymbol{\alpha}\cdot\boldsymbol{u}|)\},\quad\boldsymbol{u}\in\mathbb{S}^{1}.$
By Pythagoras, we have, for any $\boldsymbol{u}\in\mathbb{S}^{1}$,
$|\boldsymbol{\alpha}\cdot\boldsymbol{u}|^{2}+|T\boldsymbol{\alpha}\cdot\boldsymbol{u}|^{2}=|\boldsymbol{\beta}\cdot\boldsymbol{u}|^{2}+|T\boldsymbol{\beta}\cdot\boldsymbol{u}|^{2}=\|\boldsymbol{u}\|_{2}^{2}=1.$
It follows that
$|\boldsymbol{\alpha}\cdot\boldsymbol{u}|^{2}-|T\boldsymbol{\beta}\cdot\boldsymbol{u}|^{2}=|\boldsymbol{\beta}\cdot\boldsymbol{u}|^{2}-|T\boldsymbol{\alpha}\cdot\boldsymbol{u}|^{2},$
which in turn implies that $|\boldsymbol{\alpha}\cdot\boldsymbol{u}|\geq|T\boldsymbol{\beta}\cdot\boldsymbol{u}|~{}\Leftrightarrow~{}|\boldsymbol{\beta}\cdot\boldsymbol{u}|\geq|T\boldsymbol{\alpha}\cdot\boldsymbol{u}|$. Since $f$ is nonincreasing and $g$ is nondecreasing, we have $H(\boldsymbol{u})\leq 0$ for all $\boldsymbol{u}\in\mathbb{S}^{1}$, and thus
$\mathsf{E}[H(\boldsymbol{U})]\leq 0.$
On the other hand, it follows from the rotational symmetry of $\boldsymbol{U}$ and $T\boldsymbol{\beta}\cdot T\boldsymbol{\alpha}=\boldsymbol{\alpha}\cdot\boldsymbol{\beta}$ that
$\mathsf{E}[f(|\boldsymbol{\alpha}\cdot\boldsymbol{U})g(|\boldsymbol{\beta}\cdot\boldsymbol{U})]=\mathsf{E}[f(|T\boldsymbol{\beta}\cdot\boldsymbol{U})g(|T\boldsymbol{\alpha}\cdot\boldsymbol{U})].$
Similarly, since $\boldsymbol{\alpha}\cdot T\boldsymbol{\alpha}=T\boldsymbol{\beta}\cdot\boldsymbol{\beta}=\boldsymbol{e}_{1}\cdot\boldsymbol{e}_{2}=0$, we have
$\mathsf{E}[f(|\boldsymbol{\alpha}\cdot\boldsymbol{U})g(|T\boldsymbol{\alpha}\cdot\boldsymbol{U})]$ $=\mathsf{E}[f(|T\boldsymbol{\beta}\cdot\boldsymbol{U})g(|\boldsymbol{\beta}\cdot\boldsymbol{U})]$
$=\mathsf{E}[f(|\boldsymbol{e}_{1}\cdot\boldsymbol{U})g(|\boldsymbol{e}_{2}\cdot\boldsymbol{U})].$
Putting the last three equations together yields
$2\,\mathsf{E}[f(|\boldsymbol{\alpha}\cdot\boldsymbol{U})g(|\boldsymbol{\beta}\cdot\boldsymbol{U})]-2\,\mathsf{E}[f(|\boldsymbol{e}_{1}\cdot\boldsymbol{U})g(|\boldsymbol{e}_{2}\cdot\boldsymbol{U})]\leq 0,$
which proves (8). This concludes the proof. ∎
Appendix Appendix A Technical lemmas
The first lemma contains well-known formulas from matrix analysis for the inverse of a $2\times 2$ block matrix; see, e.g., Theorem 2.1 of Lu & Shiou, (2002).
Lemma 1.
Let $\Sigma$ and $M$ be symmetric matrices of width at least $2$, each partitioned into $2\times 2$ block matrices.
  • (i)
    If $A$ is invertible and the Schur complement $\Sigma/A=C-B^{\top}A^{-1}B$ is invertible, then
    $\Sigma=\begin{bmatrix}A&B\\ B^{\top}&C\end{bmatrix}\quad\Rightarrow\quad\Sigma^{-1}=\begin{bmatrix}A^{-1}+A^{-1}B(\Sigma/A)^{-1}B^{\top}A^{-1}&-A^{-1}B(\Sigma/A)^{-1}\\[2.84526pt] -(\Sigma/A)^{-1}B^{\top}A^{-1}&(\Sigma/A)^{-1}\end{bmatrix}.$
  • (ii)
    If $R$ is invertible and the Schur complement $M/R=P-QR^{-1}Q^{\top}$ is invertible, then
    $M=\begin{bmatrix}P&Q\\ Q^{\top}&R\end{bmatrix}\quad\Rightarrow\quad M^{-1}=\begin{bmatrix}(M/R)^{-1}&-(M/R)^{-1}QR^{-1}\\[2.84526pt] -R^{-1}Q^{\top}(M/R)^{-1}&R^{-1}+R^{-1}Q^{\top}(M/R)^{-1}QR^{-1}\end{bmatrix}.$
The second lemma states that multivariate normal random vectors follow the convex order according to the Loewner order of their covariance matrices; see, e.g., Shaked & Shanthikumar, (2007, Example 7.A.13).
Lemma 2.
Let $\Sigma_{1},\Sigma_{2}$ be two $d\times d$ positive definite matrices, and suppose that $\boldsymbol{Y}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma_{1})$ and $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma_{2})$. If $\Sigma_{1}\preceq\Sigma_{2}$, then $\boldsymbol{Y}$ is smaller or equal to $\boldsymbol{X}$ in the convex order, written $\boldsymbol{Y}\leq_{\mathrm{cx}}\boldsymbol{X}$, meaning that for every convex function $\psi:\mathbb{R}^{d}\to\mathbb{R}$, we have
$\mathsf{E}[\psi(\boldsymbol{Y})]\leq\mathsf{E}[\psi(\boldsymbol{X})].$
Funding
Guolie Lan’s research is supported by the National Natural Science Foundation of China (No. 12171335) and the Guangdong Provincial Natural Science Foundation (No. 2023A1515012170). Frédéric Ouimet’s previous Research Associate position at McGill University was funded through Christian Genest’s research grants (Grant RGPIN-2024-04088 [Natural Sciences and Engineering Research Council of Canada] and Grant 950-231937 [Canada Research Chairs Program]). Frédéric Ouimet’s previous postdoctoral fellowship was funded through the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2024-05794 to Anne MacKay). Wei Sun’s research is funded through the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2025-06779).